# Scheduled Curiosity-Deep Dyna-Q: Efficient Exploration for Dialog Policy Learning

### v1

@@ 3.2 Task Classifier
Turn 0 usr:
I want tickets for race at 10:00 pm.
Turn 1 sys:
What movie are you interested in?
Turn 2 usr:
I want to watch race.
Turn 3 sys:
What time would you like to see it?
Turn 4 usr:
I want to watch at 10:00 pm.
Turn 5 sys:
Which city would you like?
Turn 6 usr:
I want to watch at Seattle.
Turn 7 sys:
What date would you like to watch it?
Turn 8 usr:
I want to set it up tomorrow
Turn 9 sys:
Which theater would you like?
Turn 10 usr:
I want to watch at amc pacific place 11 theater.
Turn 11 sys:
How many tickets do you need?
Turn 12 usr:
I want 2 tickets please!
Turn 13 sys:
Great - I was able to purchase 2 tickets for you to see race tomorrow at amc pacific place 11 theater in Seattle at 10:00 pm.
Turn 14 usr:
Thank you.
Turn 15 sys:
Thank you.
@@
Intent
request, inform, deny, confirm_question,
@@
Slot
city, closing, date, distanceconstraints,
@@ 3.3 Curiosity Policy Model
Subsequently, the agent receives a reward $r_{t}$ from the environment, observes the next user action $a^{u}$, and updates to the next state $s_{t+1}$ until it reaches the end of the dialog. Experience $(s_{t},\ a_{t},\ r_{t},\ a^{u},\ s_{t+1})$ is stored in the real replay buffer $D^{u}$. We improve the Q value function $Q\left(s,\ a;\theta_{Q}\right)$ using real experiences from $D^{u}$ via minibatch SGD. It should be noted that the weight of the curiosity value cannot be adjusted for the time being. In other words, curiosity-based exploration cannot be adjusted adaptively in this study. In the planning and world model learning process, as described in the related work section, the world model is applied to interact with the agent and generate simulated experiences, which can also be used to improve the dialog policy. The planning process was similar to that of direct reinforcement learning. In this process, the environment is acted by the world model accomplished by an MLP $M\left(s,\ a;\theta_{M}\right)$ as shown in Fig. 2. In each planning turn, the world model takes the current state $s_{t}$ and the agent action $a_{t}$ as inputs. The outputs are user action $a^{u}$, reward $r_{t}$, and a binary variable $t$ indicating whether this episode is over or not. Dialog experiences generated during planning are stored in the simulated replay buffer $D^{u}$. The proposed framework follows the DDQ structure [7], which has two experience replay buffers $D^{u}$ and $D^{s}$ for storing real and simulated experiences, respectively. Because the role of the world model is to imitate real users, real experiences from $D^{u}$ were employed for its optimization via minibatch SGD. In the curiosity model training process, the curiosity model accomplished by MLP $C(s,\ a,\ \theta_{C})$ is refined via minibatch SGD using experiences from both real and simulated replay buffers, namely $D^{u}$ and $D^{s}$. Its inputs are the current state $s_{t}$ and candidate actions $a^{\prime}$, whereas its outputs are the predicted next state $s_{t+1}^{\prime}$ and curiosity value $c_{a^{\prime}}$. The labels of the predicted next state $s_{t+1}^{\prime}$ and the curiosity value $c_{a^{\prime}}$ are the real next state $s_{t+1}$ and state prediction error, respectively. To define the error of state prediction, we encode real and predicted next states $s_{t+1}$ and $s_{t+1}^{\prime}$ into vectors $\phi(s_{t+1})$ and $\phi(s_{t+1}^{\prime})$ accomplished by one-hot encoding. The state prediction error is expressed as $\ \left\|\phi\left(s_{t+1}\right)-\ \phi\left(s_{t+1}^{\prime}\right)\right\|^{2}$.
@@ 5.2 The Effectiveness of Two Opposite Training Strategies
Both typical CL and its reverse version (difficult-first strategy, DFS) benefit from the dialog policy without or with curiosity, respectively. Fig. 8 shows the success rates for each training stage. The results for the same condition were averaged. For example, the curiosity_DFS line is the average of the SC-DDQ with DME, DDE, and DDM schedules. Overall, schedules designed by the typical CL (EFS) work better on agents with the curiosity policy, and DFS is beneficial for those without the curiosity policy. Because the training of agents started with RBS, where the same rule-based policy designed for dealing with easy tasks was adopted, agents exposed to easy tasks initially showed a rapid performance boost, and the corresponding success rate reached 1.00. As the difficulty level of subsequent tasks increased, success rates decreased. Specifically, when the model without the curiosity policy attempted to grasp difficult samples after exposure to easy ones, its success rate decreased drastically, leading to poor training results in the final phase. For the model with the curiosity policy, following EFS, its success rate decreased slightly in the middle of training; however, it picked up considerably towards the end of training thanks to the encouragement of state exploration from the curiosity model.
@@
"request_slots": {
"ticket": "UNK"
"diaact": "request",
"inform slots": {
"city": "Birmingham",
"number of people": "4",
"theater": "Carmike summit 16",
"state": "al",
"start time": "around 6pm",
t "date": "today",
"movie name": "deadpool"
@@
Turn 0 usr:
Can I get 4 tickets for deadpool?
Turn 1 sys:
Which city would you like?
Turn 2 usr:
I want to watch at Birmingham.
Turn 3 sys:
Carmike summit 16 is available.
Turn 4 usr:
I want to watch at al.
Turn 5 sys:
What time would you like to see it?
Turn 6 usr:
I want to watch at around 6pm.
Turn 7 sys:
What date would you like to watch it?
Turn 8 usr:
I want to set it up today.
Turn 9 sys:
Great, I was able to purchase 4 tickets to deadpool for today at around 6pm at Carmike summit 16 in Birmingham, al.
Turn 10 usr:
Thank you.
@@
Turn 0 usr:
When is star wars playing in Oregon?
Turn 1 sys:
Portland is available.
Turn 2 usr:
What date is available?
Turn 3 sys:
12:05pm is available.
Turn 4 usr:
What date is available?
Turn 5 sys:
Friday is available.
Turn 6 usr:
Which theater is available?
Turn 7 sys:
Regal Lloyd center 10 & IMAX is available.
Turn 8 usr:
I want 4 tickets please!
Turn 9 sys:
Great, I was able to purchase 4 tickets to star wars for Friday at 12:05pm at regal Lloyd center 10 & IMAX in Portland, Oregon.
Turn 10 usr:
Thank you.
@@ 5.3 Observation of Entropy of Action Sampling
From the results shown in the previous section, we obtained a high success rate under two conditions: S-DDQ with DFS and SC-DDQ with EFS. To investigate why these conditions gave good results and other conditions did not, we introduced the entropy of sampled actions, which is used to depict the behavioral characteristics of agents. Agent sampling action preferences and entropy calculations are explained in the following paragraph. When training at different task levels, the agent prefers to sample different actions. when facing easy tasks, as shown in Fig. 9 (a), the agent tends to ask more questions instead of providing information. When facing difficult ones, as shown in Fig. 9 (b), the actions sampled by the agent tend to diversify. In other words, the agent attempts to provide appropriate information to answer the user query instead of simply asking the user questions. Furthermore, the order in which the agent meets tasks of varying difficulty affects action sampling. Therefore, to explore the effect of training schedules on agent performance, we counted the actions employed by the agent in each training phase and calculated the frequency with which each action was sampled. For example, Fig. 10 (a), (b), (c), and (d) show the frequency of actions sampled by the agent without curiosity equipped with schedule EED. On schedule EED, the agent is sequentially trained with easy, easy, and difficult tasks for 70 epochs per training stage. Finally, it is trained with all tasks for 90 epochs. The vertical coordinate is the frequency and the horizontal coordinate represents the 29 possible actions. The entropy of the distribution of sampled actions is formulated as
@@
Here, $a_{i}$ represents feasible actions. $P(a_{i})$ is the probability of choosing action $a_{i}$ during the current training stage. Entropy represents the degree of uniformity in sampling. A larger entropy implies that the agent samples actions more evenly and attempts to perform various actions rather than being trapped in a few pseudo-optimal actions. Table 6 lists the entropy values at each stage for all training conditions. Roughly speaking, in the S-DDQ with EFS conditions, the entropy values were small in the first stage and gradually increased in the following stages. The entropy values of S-DDQ with DFS showed an opposite trend. In addition, SC-DDQ with EFS showed large entropy in the first stage, whereas that with DFS showed a random up-down trend.
@@ 6 Discussion
Unlike traditional CL, in the present study, we observed that EFS and DFS had different effects on task completion ability in different task-oriented dialog models. DFS is beneficial for S-DDQ, whereas the classic CL is helpful for SC-DDQ. These two combinations share high entropy in the first stage and low entropy in the last stage, where sufficient action exploration leads to better performance. Contrary to intuition, employing a curiosity model designed to explore states is not always effective for exploring actions. To investigate the effects of the final success rate, we calculated Pearson’s correlation coefficient between the entropy values at each stage and the final success rate. The results are shown in Fig. 11. This result shows that entropy in the first stage positively correlates with the success rate, whereas entropy in the final stage is negatively correlated with the success rate. This means that an agent with a high success rate samples the actions more randomly first, and the sampling diversity converges in the last stage. If curiosity is not employed, training with easy tasks leads to a greater sampling bias and smaller entropy. This situation can be avoided using difficult tasks or curiosity. However, if we use difficult tasks at an early stage when using curiosity, the action sampling model does not converge and the success rate does not improve.
@@ 7 Limitations and Future works
This research has several limitations. First, the training schedules offered in this research were manually designed. Such schedules lack flexibility, and once designed, changes can not be made during the training process. Second, the curiosity reward cannot be adaptively adjusted for the time being. It participates in agent optimization with the same weight from beginning to end. However, dynamic adjustment of curiosity is necessary for its application in different scenarios. Finally, in the early training stage, the curiosity reward corresponding to each feasible action may deviate from its actual value. Compared with ICM [11], our curiosity model outputs the curiosity reward directly before updating to the next state, which allows the agent to obtain curiosity rewards before executing the action. Because the agent does not actually execute feasible actions individually and observe the next state, it does not obtain the real next state before acquiring the curiosity reward, and the curiosity reward output by the curiosity model is a simulation. By analyzing the entropy of action sampling, we found that the dialog system exhibits better performance when the entropy tends to decrease. In other words, we advocate that the RL-based dialog agent could benefit from encouraging exploration in the early training stage and then gradually decreasing the level of exploration. Based on our findings, we plan to add weight to the curiosity reward to make curiosity-driven exploration more adaptive. Further work on the gradual discarding of curiosity is worth conducting, because this is a generic promising path for other RL problems.
@@ 8 Conclusion
This paper presents a new framework, Scheduled Curiosity-DDQ (SC-DDQ), for task-oriented dialog policy learning. With the introduction of a curiosity model and scheduled training strategy, SC-DDQ outperforms previous classic dialog agents, such as DDQ and DQN. Moreover, variants of SC-DDQ were conducted to verify the effectiveness and influence of a typical CL and its reverse version. Based on the experimental results, the difficult-first strategy benefits S-DDQ, whereas the easy-first strategy is preferable for SC-DDQ. To explore the impact of task difficulty on the dialog agent policy, we calculated entropy. We found common trends of entropy, where the agent tried various actions randomly in the first stage, and converged in the last stage. In the future, we plan to control the curiosity-driven exploration during the training process based on entropy trends. Furthermore, we believe that this curriculum framework can be applied to improve other RL-based NLP models.

### v2

@@ 3.2 Task Classifier
Turn 0 usr: I want tickets for race at 10:00 pm.
Turn 1 sys: What movie are you interested in?
Turn 2 usr: I want to watch race.
Turn 3 sys: What time would you like to see it?
Turn 4 usr: I want to watch at 10:00 pm.
Turn 5 sys: Which city would you like?
Turn 6 usr: I want to watch at Seattle.
Turn 7 sys: What date would you like to watch it?
Turn 8 usr: I want to set it up tomorrow
Turn 9 sys: Which theater would you like?
Turn 10 usr: I want to watch at amc pacific place 11 theater.
Turn 11 sys: How many tickets do you need?
Turn 12 usr: I want 2 tickets please!
Turn 13 sys: Great - I was able to purchase 2 tickets for you to see race tomorrow at amc pacific place 11 theater in Seattle at 10:00 pm.
Turn 14 usr: Thank you.
Turn 15 sys: Thank you.
@@
Intent request, inform, deny, confirm_question,
@@
Slot city, closing, date, distanceconstraints,
@@
In this research, we designed an offline classifier focusing on booking movie tickets, identifying user goals into three complexity levels: easy, middle, and difficult. A user goal comprises two parts: request and inform slots. Inform slots are constraints known or determined by the user. The request slots are unknown to the user and must be obtained from the responses of the agent. Table 2 lists all slots. Consider Fig. 5 (b) as an example. This user goal contains three request slots: ticket, theater, and start time, and three inform slots: number of people, date, and movie name. This goal reveals that the user would like to buy three tickets for Zootopia tonight, but he/she has yet to decide which theater to go to and at a specific time. During this episode of dialog, the agent should provide information about a suitable theater and start time to satisfy the user’s constraints. In addition, the ticket is a default slot that always appears in the request slots for user goals.
Before launching the iterative scheduled SC-DDQ, the system is opened up with a warm starting phase, where a rule-based dialog policy [7] is used to interact with users and generate experiences. As the dialog strategy of the rule-based agent requires the movie name, start time, city, date, theater, and number of people in order, a user goal with a limited number of request slots is a breeze for an agent to achieve, as a successful example shown in Fig. 4. Therefore, this research applies the number of request slots as the measurement of the difficulty of training tasks, rather than the length of dialog, word rarity, and the total number of requests and inform slots [34, 13, 35, 36, 18]. The goal classification was performed offline. In this study, we classified user goals by the number of request slots, as listed in lines 2-12 of Algorithm 1.
Fig. 5 shows additional examples of the user goals at different levels. As shown in Fig. 5(a), goals with only one request slot are classified as easy and stored in the easy goal buffer $G_{\textnormal{easy}}$. Goals with two or three slots are then classified as middle and stored in the middle goal buffer $G_{\textnormal{middle}}$. Finally, goals with four or five slots are classified as difficult and stored in the difficult goal buffer $G_{\textnormal{difficult}}$. At the beginning of the dialog, the user samples a goal from the corresponding set. In this research, schedules designed on two opposite training strategies are conducted, which are Easy-Middle-Difficult (EMD), Easy-Difficult-Difficult (EDD), Easy-Easy-Difficult-All (EED), Difficult-Middle-Easy (DME), Difficult-Easy-Easy (DEE) and Difficult-Difficult-Middle (DDM).
@@ 3.3 Curiosity Policy Model
Subsequently, the agent receives a reward $r_{t}$ from the environment, observes the next user action $a^{u}$, and updates to the next state $s_{t+1}$ until it reaches the end of the dialog. Experience $(s_{t},\ a_{t},\ r_{t},\ a^{u},\ s_{t+1})$ is stored in the real replay buffer $D^{u}$. We improve the Q value function $Q\left(s,\ a;\theta_{Q}\right)$ using real experiences from $D^{u}$ via minibatch SGD. It should be noted that the weight of the curiosity value cannot be adjusted for the time being. In other words, curiosity-based exploration cannot be adjusted adaptively in this study.
In the planning and world model learning process, as described in the related work section, the world model is applied to interact with the agent and generate simulated experiences, which can also be used to improve the dialog policy. The planning process was similar to that of direct reinforcement learning. In this process, the environment is acted by the world model accomplished by an MLP $M\left(s,\ a;\theta_{M}\right)$ as shown in Fig. 2. In each planning turn, the world model takes the current state $s_{t}$ and the agent action $a_{t}$ as inputs. The outputs are user action $a^{u}$, reward $r_{t}$, and a binary variable $t$ indicating whether this episode is over or not. Dialog experiences generated during planning are stored in the simulated replay buffer $D^{u}$. The proposed framework follows the DDQ structure [7], which has two experience replay buffers $D^{u}$ and $D^{s}$ for storing real and simulated experiences, respectively. Because the role of the world model is to imitate real users, real experiences from $D^{u}$ were employed for its optimization via minibatch SGD.
In the curiosity model training process, the curiosity model accomplished by MLP $C(s,\ a,\ \theta_{C})$ is refined via minibatch SGD using experiences from both real and simulated replay buffers, namely $D^{u}$ and $D^{s}$. Its inputs are the current state $s_{t}$ and candidate actions $a^{\prime}$, whereas its outputs are the predicted next state $s_{t+1}^{\prime}$ and curiosity value $c_{a^{\prime}}$. The labels of the predicted next state $s_{t+1}^{\prime}$ and the curiosity value $c_{a^{\prime}}$ are the real next state $s_{t+1}$ and state prediction error, respectively. To define the error of state prediction, we encode real and predicted next states $s_{t+1}$ and $s_{t+1}^{\prime}$ into vectors $\phi(s_{t+1})$ and $\phi(s_{t+1}^{\prime})$ accomplished by one-hot encoding. The state prediction error is expressed as $\ \left\|\phi\left(s_{t+1}\right)-\ \phi\left(s_{t+1}^{\prime}\right)\right\|^{2}$.
@@ 5.2 The Effectiveness of Two Opposite Training Strategies
Both typical CL and its reverse version (difficult-first strategy, DFS) benefit from the dialog policy without or with curiosity, respectively. Fig. 8 shows the success rates for each training stage. The results for the same condition were averaged. For example, the curiosity_DFS line is the average of the SC-DDQ with DME, DDE, and DDM schedules. Overall, schedules designed by the typical CL (EFS) work better on agents with the curiosity policy, and DFS is beneficial for those without the curiosity policy.
Because the training of agents started with RBS, where the same rule-based policy designed for dealing with easy tasks was adopted, agents exposed to easy tasks initially showed a rapid performance boost, and the corresponding success rate reached 1.00. As the difficulty level of subsequent tasks increased, success rates decreased. Specifically, when the model without the curiosity policy attempted to grasp difficult samples after exposure to easy ones, its success rate decreased drastically, leading to poor training results in the final phase. For the model with the curiosity policy, following EFS, its success rate decreased slightly in the middle of training; however, it picked up considerably towards the end of training thanks to the encouragement of state exploration from the curiosity model.
@@
{ "request_slots": { "ticket": "UNK" }, "diaact": "request", "inform slots": { "city": "Birmingham", "number of people": "4", "theater": "Carmike summit 16", "state": "al", "start time": "around 6pm", t "date": "today", "movie name": "deadpool" } }
@@
Turn 0 usr: Can I get 4 tickets for deadpool?
Turn 1 sys: Which city would you like?
Turn 2 usr: I want to watch at Birmingham.
Turn 3 sys: Carmike summit 16 is available.
Turn 4 usr: I want to watch at al.
Turn 5 sys: What time would you like to see it?
Turn 6 usr: I want to watch at around 6pm.
Turn 7 sys: What date would you like to watch it?
Turn 8 usr: I want to set it up today.
Turn 9 sys: Great, I was able to purchase 4 tickets to deadpool for today at around 6pm at Carmike summit 16 in Birmingham, al.
Turn 10 usr: Thank you.
@@
Turn 0 usr: When is star wars playing in Oregon?
Turn 1 sys: Portland is available.
Turn 2 usr: What date is available?
Turn 3 sys: 12:05pm is available.
Turn 4 usr: What date is available?
Turn 5 sys: Friday is available.
Turn 6 usr: Which theater is available?
Turn 7 sys: Regal Lloyd center 10 & IMAX is available.
Turn 8 usr: I want 4 tickets please!
Turn 9 sys: Great, I was able to purchase 4 tickets to star wars for Friday at 12:05pm at regal Lloyd center 10 & IMAX in Portland, Oregon.
Turn 10 usr: Thank you.
@@ 5.3 Observation of Entropy of Action Sampling
From the results shown in the previous section, we obtained a high success rate under two conditions: S-DDQ with DFS and SC-DDQ with EFS. To investigate why these conditions gave good results and other conditions did not, we introduced the entropy of sampled actions, which is used to depict the behavioral characteristics of agents. Agent sampling action preferences and entropy calculations are explained in the following paragraph.
When training at different task levels, the agent prefers to sample different actions. when facing easy tasks, as shown in Fig. 9 (a), the agent tends to ask more questions instead of providing information. When facing difficult ones, as shown in Fig. 9 (b), the actions sampled by the agent tend to diversify. In other words, the agent attempts to provide appropriate information to answer the user query instead of simply asking the user questions.
Furthermore, the order in which the agent meets tasks of varying difficulty affects action sampling. Therefore, to explore the effect of training schedules on agent performance, we counted the actions employed by the agent in each training phase and calculated the frequency with which each action was sampled. For example, Fig. 10 (a), (b), (c), and (d) show the frequency of actions sampled by the agent without curiosity equipped with schedule EED. On schedule EED, the agent is sequentially trained with easy, easy, and difficult tasks for 70 epochs per training stage. Finally, it is trained with all tasks for 90 epochs. The vertical coordinate is the frequency and the horizontal coordinate represents the 29 possible actions. The entropy of the distribution of sampled actions is formulated as
@@
Here, $a_{i}$ represents feasible actions. $P(a_{i})$ is the probability of choosing action $a_{i}$ during the current training stage. Entropy represents the degree of uniformity in sampling. A larger entropy implies that the agent samples actions more evenly and attempts to perform various actions rather than being trapped in a few pseudo-optimal actions.
Table 6 lists the entropy values at each stage for all training conditions. Roughly speaking, in the S-DDQ with EFS conditions, the entropy values were small in the first stage and gradually increased in the following stages. The entropy values of S-DDQ with DFS showed an opposite trend. In addition, SC-DDQ with EFS showed large entropy in the first stage, whereas that with DFS showed a random up-down trend.
@@ 6 Discussion
Unlike traditional CL, in the present study, we observed that EFS and DFS had different effects on task completion ability in different task-oriented dialog models. DFS is beneficial for S-DDQ, whereas the classic CL is helpful for SC-DDQ. These two combinations share high entropy in the first stage and low entropy in the last stage, where sufficient action exploration leads to better performance. Contrary to intuition, employing a curiosity model designed to explore states is not always effective for exploring actions.
To investigate the effects of the final success rate, we calculated Pearson’s correlation coefficient between the entropy values at each stage and the final success rate. The results are shown in Fig. 11. This result shows that entropy in the first stage positively correlates with the success rate, whereas entropy in the final stage is negatively correlated with the success rate. This means that an agent with a high success rate samples the actions more randomly first, and the sampling diversity converges in the last stage. If curiosity is not employed, training with easy tasks leads to a greater sampling bias and smaller entropy. This situation can be avoided using difficult tasks or curiosity. However, if we use difficult tasks at an early stage when using curiosity, the action sampling model does not converge and the success rate does not improve.
@@ 7 Limitations and Future works
This research has several limitations. First, the training schedules offered in this research were manually designed. Such schedules lack flexibility, and once designed, changes can not be made during the training process. Second, the curiosity reward cannot be adaptively adjusted for the time being. It participates in agent optimization with the same weight from beginning to end. However, dynamic adjustment of curiosity is necessary for its application in different scenarios. Finally, in the early training stage, the curiosity reward corresponding to each feasible action may deviate from its actual value. Compared with ICM [11], our curiosity model outputs the curiosity reward directly before updating to the next state, which allows the agent to obtain curiosity rewards before executing the action. Because the agent does not actually execute feasible actions individually and observe the next state, it does not obtain the real next state before acquiring the curiosity reward, and the curiosity reward output by the curiosity model is a simulation.
By analyzing the entropy of action sampling, we found that the dialog system exhibits better performance when the entropy tends to decrease. In other words, we advocate that the RL-based dialog agent could benefit from encouraging exploration in the early training stage and then gradually decreasing the level of exploration. Based on our findings, we plan to add weight to the curiosity reward to make curiosity-driven exploration more adaptive. Further work on the gradual discarding of curiosity is worth conducting, because this is a generic promising path for other RL problems.
@@ 8 Conclusion
This paper presents a new framework, Scheduled Curiosity-DDQ (SC-DDQ), for task-oriented dialog policy learning. With the introduction of a curiosity model and scheduled training strategy, SC-DDQ outperforms previous classic dialog agents, such as DDQ and DQN. Moreover, variants of SC-DDQ were conducted to verify the effectiveness and influence of a typical CL and its reverse version. Based on the experimental results, the difficult-first strategy benefits S-DDQ, whereas the easy-first strategy is preferable for SC-DDQ. To explore the impact of task difficulty on the dialog agent policy, we calculated entropy. We found common trends of entropy, where the agent tried various actions randomly in the first stage, and converged in the last stage.
In the future, we plan to control the curiosity-driven exploration during the training process based on entropy trends. Furthermore, we believe that this curriculum framework can be applied to improve other RL-based NLP models.